OpenStudy (anonymous):

Solve the system by substitution. y=x*2-3 y*2=4

OpenStudy (anonymous):

Find the value of Y: y*2=4

OpenStudy (anonymous):

should be y=4/2

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

then plug in that 2 in for the Y in y=x*2-3

OpenStudy (anonymous):

and solve for X :)